# A Mapping Analysis of Requirements Between the CRA and the GDPR

### v1

@@ Abstract
Index Terms:
Legal requirements, essential requirements, cyber security, regulations, compliance, conformance, redundancy
@@ II Related Work
Due to other recent ripples in the EU’s legal framework for cyber security, the GDPR has also been considered in conjunction with other recent cyber security laws [28]. Thus far, however, existing research on the CRA is scant but not entirely absent, the likely explanation being simple; at the time of writing, the regulation is new and the details are still being worked on in the member states. Regarding research, besides legal overviews [29], the CRA’s new requirements for vulnerability coordination and disclosure have been recently addressed [43]. There is also work on the CRA’s still unclear implications upon open source software [4]. The regulation’s requirement about secure default configurations has also been implicitly addressed through a literature review [38]. These and potentially other outlying examples notwithstanding, the volume of existing research is still low. For this reason, the CRA also takes a priority in the mapping in the form of Q.2. Further related work can be pointed out by considering some of the core concepts together with additional framings.
@@ III-B Requirement Types
A further clarification is required: the paper operates with legal requirements, which are generally about something that a law requires for software, systems, and other implementations. The term legal requirement is important for three reasons. First, it separates law-imposed requirements from “conventional” requirements, such as those elicited from clients and customers. With respect to non-functional requirements in general, legal requirements are particularly something that are non-negotiable [47]. Second, compliance with legal requirements can be seen as an overall non-functional requirement from which, if possible, functional requirements should be drawn. As non-functional requirements too should be verifiable [47], this point reiterates the point about processes. Third, legal requirements are also a distinct genre within requirements engineering research. Despite various models, frameworks, techniques, and tools, the challenges in this genre are well-recognized [15, 33]. In what follows, no specific methodology is required because the GDPR’s requirements have already been addressed in previous research, and the CRA’s requirements are relatively straightforward to identify (but not necessarily interpret) because the noted framing toward the regulation’s essential requirements. This term needs a clarification too. It also serves to pinpoint a few brief points about compliance and conformance with the CRA regulation.
@@ V-C Security Updates
Even though optional, automatic updates can be seen as particularly relevant for improving the IoT domain’s cyber security posture [3]. The provisions and options provided also indicate that the CRA seems to have taken into account most—or even all—of the deadly sins associated with software updates, as elaborated in the literature [17]. However, the CRA’s wide scope, which includes everything from smartphones to industrial control systems, will likely bring new challenges for some vendors. In addition, there are always risks involved too in updating; in fact, annual surveys indicate that faulty updates have been relatively common in some industry sectors [10]. Regressions are also unfortunately common in many software projects, including critical ones, such as the Linux kernel [39]. These and other potential problems warrant also further research in software engineering and beyond.
@@ VI-B Further Overlaps
As has been pointed out [29], there is a further overlap between the two regulations and the radio equipment Directive 2014/53/EU. There are thee obligations specified in the directive’s Article 3(3)(d), (e), and (f) that overlap with the CRA and the GDPR: personal data should be protected, fraud prevention should be implemented when possible, and an “equipment does not harm the network or its functioning nor misuse network resources, thereby causing an unacceptable degradation of service”. The obligation quoted resembles the point about cascading effects noted in Subsection IV-A.
As has again been pointed out [41], the GDPR also overlaps slightly with the EU’s recent Directive (EU) 2022/2555, which is commonly known as the NIS2 directive. The overlap comes from incident notifications about serious personal data breaches to public authorities. Although a direct overlap is not present, also the CRA imposes incident notification obligations. The regulation’s Article 14 mandates manufacturers to report about actively exploited vulnerabilities as well as severe incidents to the supervising authorities. While the deadlines of reporting align with those specified in the NIS2 directive, the definitions for severe incidents are different. According to the CRA’s Article 14(5), these refer to those that affect the CIA guarantees of products on one hand, and those “leading to the introduction or execution of malicious code in a product with digital elements or in the network and information systems of a user of the product” on the other hand. Once again, the theme of cascading effects is present in this direct quotation.
@@ VI-C Further Work
Further work is required also on risk analysis and auditing. While the details are still being worked on in the EU, a good topic related to risk analysis would be an evaluation of the suitability of existing standards for auditing, whether internally or externally, a fulfillment of the CRA’s essential requirements. Although there is existing work in this area [11], risk analysis itself seems to have received less attention. As was noted in Subsection III-D, risk analysis should guide the prioritization of the essential requirements and their implementations.
Auditing would be a good topic also otherwise because cyber security audits are only rarely conducted according to some industry surveys [35]. Regarding the survey results referenced, it seems that the CRA patches also some other notable limitations in industry practices. Although costs are certainly a factor, also attack surface analysis and security reviews more generally seem to be only occasionally done in the industry, among other things. In light of such observations, another industry survey about the CRA specifically would also shed light on the potential challenges ahead regarding compliance. As already remarked, it be that the challenges are sectoral, specific to some particular products, or both. For instance, industrial control systems and supervisory control and data acquisition (SCADA) systems have been seen to suffer from many of the issues addressed by the CRA. That is, among other things, these systems do not always have robust authentication and authorization methods, remote access contains particular risks, updates are often lacking, and monitoring and logging are oftentimes insufficient [1]. These and other real or perceived challenges correlate with real or perceived problems in addressing them. For instance: due to the criticality of the systems in many industry sectors, including those related to countries’ critical infrastructures, integrity guarantees and pre-delivery and post-delivery verification of updates are likely of significant importance. That is, faulty updates alone are likely a high-impact risk with these systems and in these sectors.

### v2

@@ II Related Work
Due to other recent ripples in the EU’s legal framework for cyber security, the GDPR has also been considered in conjunction with other recent cyber security laws [28]. Thus far, however, existing research on the CRA is scant but not entirely absent, the likely explanation being simple; at the time of writing, the regulation is new and the details are still being worked on in the member states. Regarding research—and in addition to legal overviews [29], the CRA’s new obligations for vulnerability coordination and disclosure have been discussed [46]. Recently, the alignment of the CRA’s essential requirements and the MITRE’s ATT&CK® mitigations has been evaluated [43]. There is also work on the CRA’s still unclear implications upon open source software [4]. The regulation’s requirement about secure default configurations has also been implicitly addressed through a literature review [39]. These and potentially other outlying examples notwithstanding, the volume of existing research is still low. For this reason, the CRA also takes a priority in the mapping in the form of Q.2. Further related work can be pointed out by considering some of the core concepts together with additional framings.
@@ III-B Requirement Types
A further clarification is required: the paper operates with legal requirements, which are generally about something that a law requires for software, systems, and other implementations. The term legal requirement is important for three reasons. First, it separates law-imposed requirements from “conventional” requirements, such as those elicited from clients and customers. With respect to non-functional requirements in general, legal requirements are particularly something that are non-negotiable [51]. Second, compliance with legal requirements can be seen as an overall non-functional requirement from which, if possible, functional requirements should be drawn. As non-functional requirements too should be verifiable [51], this point reiterates the point about processes. Third, legal requirements are also a distinct genre within requirements engineering research. Despite various models, frameworks, techniques, and tools, the challenges in this genre are well-recognized [15, 34]. In particular, legal requirements are methodologically challenging in requirements engineering research because these require an interdisciplinary approach. Although different ontologies and related techniques have been presented [50], elicitation of legal requirements necessitates actually reading and interpreting law, no more or no less. This truism explains also why a collaboration between engineers and lawyers is often desirable in the industry when dealing with legal requirements [15]. Reading and interpreting are also what the paper’s methodological approach for the mapping analysis is about—as is the “methodology” in legal scholarship in general. While these provide a basic toolbox to make the mappings, a step toward more traditional requirements engineering is taken by tentatively further interpreting how the legal requirements identified translate into functional requirements for software, systems, and other implementations
@@ V-C Security Updates
Even though optional, automatic updates can be seen as particularly relevant for improving the IoT domain’s cyber security posture [3]. The provisions and options provided also indicate that the CRA seems to have taken into account most—or even all—of the deadly sins associated with software updates, as elaborated in the literature [17]. However, the CRA’s wide scope, which includes everything from smartphones to industrial control systems, will likely bring new challenges for some vendors. In addition, there are always risks involved too in updating; in fact, annual surveys indicate that faulty updates have been relatively common in some industry sectors [10]. Faulty updates have also led to some memorable, severe, but non-security incidents over the years [32]. Furthermore, regressions are unfortunately common in many software projects, including critical ones, such as the Linux kernel [40]. These and other potential problems warrant also further research in software engineering and beyond.
@@ VI-B Further Overlaps
As has been pointed out [29], there is a further overlap between the two regulations and the radio equipment Directive 2014/53/EU. There are three obligations specified in the directive’s Article 3(3)(d), (e), and (f) that overlap with the CRA and the GDPR: personal data should be protected, fraud prevention should be implemented when possible, and an “equipment does not harm the network or its functioning nor misuse network resources, thereby causing an unacceptable degradation of service”. The obligation quoted resembles the point about cascading effects noted in Subsection IV-A.
As has again been pointed out [42], the GDPR also overlaps slightly with the EU’s recent Directive (EU) 2022/2555, which is commonly known as the NIS2 directive. The overlap comes from incident notifications about serious personal data breaches to public authorities. Although a direct overlap is not present, also the CRA imposes incident notification obligations. The regulation’s Article 14 mandates manufacturers to report about actively exploited vulnerabilities as well as severe incidents to the supervising authorities. While the deadlines of reporting align with those specified in the NIS2 directive [44], the definitions for severe incidents are different. According to the CRA’s Article 14(5), these refer to those that affect the CIA guarantees of products on one hand, and those “leading to the introduction or execution of malicious code in a product with digital elements or in the network and information systems of a user of the product” on the other hand. Once again, the theme of cascading effects is present in this direct quotation. As the NIS2 directive obliges operators of critical infrastructures to notify about other incident types too and conduct comprehensive risk analyses, there is also a more broad overlap with the CRA. Seeking synergies between these two cyber security laws may be worthwhile in practice.
Furthermore, there is an overlap with Regulation (EU) 2023/1670 addressing mobile phones, including smartphones. This overlap is about product life cycles, including the obligation to provide security updates discussed in Subsection V-C and the sustainability theme noted in Subsection IV-D. That is, the mobile phone regulation mandates phone manufacturers to provide operating system updates for a minimum of five years.
@@ VI-C Further Work
Further work is required also on risk analysis and auditing. While the details are still being worked on in the EU, a good topic related to risk analysis would be an evaluation of the suitability of existing standards for auditing, whether internally or externally, a fulfillment of the CRA’s essential requirements. Although there is existing work in this area [11], risk analysis itself seems to have received less attention. As was noted in Subsection III-D, risk analysis should guide the prioritization of the essential cyber security requirements and their implementations. As was also noted, however, a risk analysis requires a case or a uniform domain; universal guidelines are impossible to postulate because of the CRA’s wide scope.
Auditing would be a good topic also otherwise because cyber security audits are only rarely conducted according to some industry surveys [36]. Regarding the survey results referenced, it seems that the CRA patches also some other notable limitations in industry practices. Although costs are certainly a factor, also attack surface analysis and security reviews more generally seem to be only occasionally done in the industry [36]. In light of such observations, another industry survey about the CRA specifically would also shed light on the potential challenges ahead regarding compliance.
As already remarked, it be that the challenges are sectoral, specific to some particular products, or both. For instance, industrial control systems and supervisory control and data acquisition (SCADA) systems have been seen to suffer from many of the issues addressed by the CRA. That is, among other things, these systems do not always have robust authentication and authorization methods, remote access contains particular risks, updates are often lacking, and monitoring and logging are oftentimes insufficient [1]. These and other real or perceived challenges correlate with real or perceived problems in addressing them. For instance: due to the criticality of the systems in many industry sectors, including those related to countries’ critical infrastructures, integrity guarantees and pre-delivery and post-delivery verification of updates are likely of significant importance. That is, faulty updates alone are likely a high-impact risk with these systems and in these sectors.